Grade 12Electric Current

In H.C Vermas concepts of physics vol-2 in the page.no.106 in Example 29.1, how is the angle cosine theta determined ???

There i sno explaination whatsoever in the book plz... help me im a weak student in electricity chapters

Pls...help Pls i will approve the ans anyway help me aieee approaching n givin me the spine chills !!!

Geoffrey Richards
15 Years agoGrade 12

2 Answers

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

see the triangle APBC , which is an isoceles triangle ...

after dropping a perpendicular from P to base AB , triangle is divided into two halves ..

PBC is one of the triangle which is right angled , base of this triangle is 5 ...

now applying trigonometry ,

tan@ = perpendicular/base = 12/5

sec2@ = 1+tan2@ = 169/25 so

cos@ = 5/13

Geoffrey Richards
15 Years ago

I'm surprised !!!!

Vikas are u a iitian ???

Plsss reply if u are there!!!